OpenStudy (anonymous):

In radian measure, the angles of a 45-45-90 triangle are:

OpenStudy (anonymous):

\(\frac{\pi}{4},\frac{\pi}{4},\frac{\pi}{2}\)

OpenStudy (anonymous):

divide by \(180\) and stick a \(\pi\) next to it

OpenStudy (anonymous):

can I ask how it's done? ohhh ok I see

OpenStudy (anonymous):

\[\frac{45}{180}=\frac{1}{4}\] so \(\frac{\pi}{4}\)

OpenStudy (anonymous):

Ok that makes sense

OpenStudy (anonymous):

good !